how can i find the perimeter of a rectangle that is 70m and the length is 5m more than twice the width find the dimension.
Found 2 solutions by richwmiller, georgiabh:
Answer by richwmiller(17219)

You asked" How can I"
You can hire someone to do it for you.
You can learn how to do it on your own.
or
perimeter=70
Since you want to know the perimeter we are done.
But my guess is that you copied the problem incorrectly.
It is hard to get a correct answer with a wrong question.

L=5+2w
p=2L+2W
70=2*(5+2w)+2w
70=10+4w+2w
60=6w
10=w
L=25
length is 25 and width is 10

Answer by georgiabh(4)

The question doesn't seem quite right!
If you wish to say that the perimeter is 70m and you want to find the length of each side then you can solve as follows.
A rectangle has opposite sides of equal length.
We can call the width x.
The length is 5m more than twice the width so we say that the length is 2x + 5.
The perimeter is the whole way round the rectangle which is x + 2x + 5 + x + 2x + 5 = 70m
So, 6x + 10 = 70
6x = 60
x = 10m and 2x + 5 = 25
so the dimensions of the rectangle are 10 x 25m.